OpenStudy (anonymous):

Hi, i want to calculate analytically a double integral in polar coordinate. I need to plot the function in terms of variable r?

OpenStudy (anonymous):

Like\[g(x)= \int\limits\limits_{0}^{r}\int\limits\limits_{0}^{2\pi } f \left( \ x,\rho,\theta \right) \rho d \theta d \ \rho \]?

OpenStudy (anonymous):

Hello, yes but i need to plot it

OpenStudy (turingtest):

Once you have integrated you will have a function in terms of x. Graph whatever that is. I am going to close this since it belongs in the math section, you may repost it there.

OpenStudy (anonymous):

ok thanks